OpenStudy (anonymous):

7th grade, currently doing square roots. The question was What are the square roots of 400? The answer was The number 400 has a positive and a negative square root. Also, 400 = 20 to the second power. So the square roots of 400 are 20 and -20. Can someone explain how they came to that????

OpenStudy (anonymous):

20 times 20 is 400

OpenStudy (amorfide):

\[\sqrt{400}=\sqrt{x^{2}}\] you can rewrite the number under the root sign if you think of a square number that goes into 400, we will choose 4, and multiply it by a number to make 400 so 4 multiplied by 100 \[\sqrt{c}=\sqrt{a \times b}\] \[\sqrt{400}=\sqrt{4 \times 100}\] you can also seperate this by following the rule \[\sqrt{a \times b}=\sqrt{a} \times \sqrt{b}\] \[\sqrt{4 \times 100} = \sqrt{4} \times \sqrt{100}\] \[\sqrt{4}=2\] \[\sqrt{100}=10\] 2x10=20 and as you know, you always get a positive and negative answer since \[(-20) \times (-20)=(-1 \times -1)(20 \times 20)\] negative multiplied by negative is a positive or you can just remember the answer to the square of the first 20 numbers

OpenStudy (anonymous):

I pretty much know the square root of 1-12 or so using 2.. Plus I really suck at Pre-Algebra Thanks for the help!

OpenStudy (amorfide):

to be honest you could just do trial and error, but you should memorise the squares of a number that is a multiple of 10 10x10=100 20x20=400 30x30=900 they are pretty easy since for 10 you square 1 then put as many zeros as you see since 10x10 there are 2 zeros you get 100 for 20 2squared is 4 two zeros 400 so if you had to find the square root of 361 you know it is close to 20 squared since it is close to 400 and you know it is more than 10 squared so try finding the square of a number half way between 10 and 20 15 15 squared is 225 it is between 15 and 20 try 17 17 squared is 289 between 17 and 20 try 19 19 squared is 361 more long winded idk what way you prefer